OpenStudy (anonymous):

Will Medal :) Which equation can be used to find the two numbers whose ratio is 3 to 2 and that have a sum of 35?

OpenStudy (anonymous):

Number one = 2x Number two =3x 2x +3x = 35 5x=35 x=7 Number one is 2*7=14 Number two is 3*7=21